A long, straight wire of diameter 5.0 mm carries a uniformly distributed current of
15 A. At what distance from the axis of the wire will the magnitude of B be maximum? Justify your answer.
1
Expert's answer
2020-03-04T10:00:15-0500

"R=5.0 mm"

"I=15A"

According to Ampere's circuital law

"\int{Bdl=\mu_0I}"


If "r<R"


"B\cdot 2\pi r=\mu_0\cdot \frac{I}{4\pi R^2}\cdot 4\pi r^2\to B=\frac{\mu_0I}{2\pi R^2}\cdot r"


If "r>R"


"B\cdot 2\pi r=\mu_0\cdot I\to B=\frac{\mu_0I}{2\pi r}"


So, if "r=R=5.0 mm" then "B=B_{max}"


"B_{max}=\frac{\mu_0I}{2\pi R}=\frac{4\cdot \pi\cdot 10^{-7}\cdot15}{2\cdot\pi \cdot5\cdot 10^{-3}}=6\cdot 10^{-4}T"